Sean’s Notebook: Restarting Canu

Looks like they took Mox down for a bit last night/this morning for maintenance so our Canu run got bumped. Thankfully it can fail gracefully and restart midway so hopefully it will be done soon.

Ive got my long read samples concatenated and ready for consensus calling, next I need to organize the short read stuff by insert sizes for polishing, and see what additional prep needs to be done for them.

Laura’s Notebook: Manchester, 6/10

Maintenance day. Steven and Beyer helped me, and it was a productive morning.

Activities:

  • I counted the larvae that from yesterday’s screening, then imaged them. I put the well plates in the fridge yesterday without fixing in Lugols. I thus imaged the larvae still alive, but sluggish from the cold.
  • SR & BR cleaned lines. They ran freshwater through the whole system first for 10 minutes while using the algae dosing pump to inject a mild bleach solution (1mL 12.5% chlorine/mL fresh water), then cleansed the line with fresh water for 10 minutes.
  • SR & BR replaced all drippers, banjos, and cleaned the larval catchment buckets.
  • I collected new larvae:
    • NF-6 Low A
    • K-6 Ambient
    • K-10 Ambient (just a few, not healthy, did not stock)
    • K-6 Low (just a few, not healthy, did not stock)
    • HL-10 Low
  • Checked out things hanging off of the pier, in an attempt to find the rest of Katherine’s oysters. Candidate was tag #7 in the pearl nets. UPDATE: Confirmed with Stuart that this is where the oysters are. Need to return to the pier, open those up, and see what groups are there. This will help me confirm the group that I have in my experiment.

Took Sunday off; set up extra algae header tank on Saturday.

from The Shell Game http://ift.tt/2szemhA
via IFTTT